The length of the top of the table is 2 m greater than the width. The area is 80m^2. Find the dinensions of the table.

Answer by mananth(16946)

Area = 80 sq. ft
Length = x ft
Width ( x -2 )
Area = x * ( x -2 )
80 = x^2 + -2 x
x^2 -2 x -80 = 0
x^2-10x+8x-80=0
x(x-10)+8(x-10)=0
(x-10)(x+8)=0
x= 10 ignoring negative
dimensions are 10 & 8 ft